# Non-radial oscillations of hadronic neutron stars, quark stars, and hybrid stars : Estimation of $f$, $p$, and $g$ modes

### v1

@@ II.1.1 KIDS functional
In addition to the rules for the expansion of the energy, KIDS formalism also assumes rules for fitting the model parameters. Eq. (1) represents the energy in homogeneous infinite nuclear matter. Three $\alpha_{i}$ and four $\beta_{i}$ are adjusted to the input nuclear matter data. To apply the theory to finite nuclei, we transform the energy to Skyrme-type contact interactions so that it is easily implemented in the Hartree-Fock codes. Details for the steps toward finite nuclei are illustrated in Refs. Gil et al. (2017a, b); Papakonstantinou and Hyun (2023). In the Skyrme-type force, two terms describing the density gradient and the spin-orbit coupling are added to the nuclear matter functional. Two new parameters in the Skyrme-type force are fitted to reproduce selected nuclear data. Since the fitting to nuclear data is done on top of the parameters determined from the nuclear matter data, nuclear matter EoS can be treated independently of the nuclear properties.
@@
Following the rules described above, KIDS-A, B, C and D models are determined to satisfy the neutron star properties constrained by modern astronomical observations. Density dependence of the symmetry energy plays a crucial role in determining the nuclear matter EoS at densities below and above the saturation density. With the purpose to reduce the uncertainties in the density dependence of the symmetry energy, incompressibility of the symmetric matter $K_{0}$, and the symmetry energy parameters $J$, $L$ and $K_{\rm sym}$ are constrained to reproduce the nuclear data and neutron star observation simultaneously. Detailed process is explained in Gil and Hyun (2021). Table LABEL:Tab:KIDS tabulates the values of $K_{0}$, $J$, $L$ and $K_{\rm sym}$ for the KIDS-A, B, C and D models.
@@ II.1.2 RMF models
The nucleons interact via the scalar $\sigma$, vector $\omega$ and iso-vector $\rho$ mesons. The vacuum expectation values of the meson fields ($\sigma_{0}$, $\omega_{0}$ and $\rho_{03}$) in RMF approximation and the EoS obtained from the Lagrangian (Eq. (2)) can be found in Xia et al. (2022). For the GM1 model the mesons have density-independent couplings $g_{\sigma}$, $g_{\omega}$ and $g_{\rho}$ with the nucleons and $g_{2}$ and $g_{3}$ are the higher order scalar field coefficients while $c_{3}$ is the higher order vector field coefficient. These non-linear self couplings are effectively considered in order to account for the in-medium effects. On the other hand, in the DD2 model Typel et al. (2010) $g_{2}$=$g_{3}$=$c_{3}$=0 and the in-medium effects are treated with the density-dependent couplings $g_{i}(\rho)$ (where, $i=\sigma,\omega,\rho$) following the Typel-Wolter ansatz Lu et al. (2011). The values of the density-independent couplings for GM1 model and density-dependent couplings for DD2 model along with the other parameters of the two models can be found in the respective references and also in Guha and Sen (2024). The saturation properties like the saturation density $\rho_{0}$, symmetry energy $J$, slope $L$, nuclear incompressibility $K_{0}$, skewness coefficient $S_{0}$, and the curvature parameter $K_{\rm{sym}}$ of the nuclear symmetry energy, as obtained for the two RMF models considered in this present work, can be found in the respective references and also in Xia et al. (2022). For comparison with the KIDS model, we also display the values of $K_{0}$, $J$, $L$ and $K_{\rm sym}$ for DD2 and GM1 models in Table LABEL:Tab:KIDS.
@@ II.2.1 Vector MIT bag (v-Bag) model
We consider that the quark matter is composed of the $u$, $d$ and $s$ quarks and the electrons. The masses of the quarks are $m_{u}$ = 5 MeV, $m_{d}$ = 7 MeV and $m_{s}$ = 95 MeV. Based on the MIT bag model framework Chodos et al. (1974) , in the v-Bag model the repulsion between the quarks is mediated by a vector meson of mass $m_{V}$ = 783 MeV Lopes et al. (2022); Kumar et al. (2022); Laskos-Patkos et al. (2024); Kumar et al. (2023c). The Lagrangian for the v-Bag model is given by
@@
where, $V_{\mu}$ denotes the vector field, $g_{qqV}$ is the quark-vector meson coupling constant, $B$ is the bag constant, $V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$ and the Heaviside function $\Theta$=1 inside the bag. For the vector coupling constant, we assume $g_{uuV}=g_{ddV}$ and $X_{V}=g_{ssV}/g_{uuV}$ = 0.4. In case of QSs, the parameter $G_{V}=(g_{uuV}/m_{V})^{2}$ and its corresponding value of $B$ are constrained by the Bodmer-Witten conjecture which is related to the stability of the star. For a particular value of $G_{V}$, $B_{max}$ are determined by the binding energy of the star at the surface $\varepsilon/\rho_{B}\leq$ 930 MeV while $B_{min}$ is determined with the 2-flavor QM Farhi and Jaffe (1984); Torres and Menezes (2013); Lopes et al. (2022). The EoS for SQM derived from the Lagrangian (Eq. (3)) can be found in Lopes et al. (2022). The parameter set, used in the present work for obtaining the SQS configurations with the v-Bag model, is tabulated in Tab. (2).
@@ II.2.2 Nambu-Jona-Lasino (NJL) model
${\cal L}_{vNJL}=\bar{\psi}_{f}[\gamma^{\mu}(i\partial_{\mu}-m_{f})]{\psi}_{f}+G_{S}\sum_{a=0}^{8}\left[(\bar{\psi}_{f}\lambda_{a}\psi)^{2}+(\bar{\psi}_{f}\gamma_{5}\lambda_{a}\psi)^{2}\right]$
$-G_{V}(\bar{\psi}_{f}\gamma^{\mu}\psi)^{2}-K\left\{det[\bar{\psi}(1+\gamma_{5})\psi]+det[\bar{\psi}(1-\gamma_{5})\psi]\right\}$ (4)
where, $m_{f}$ are the current quark masses, $\lambda_{a}$ are the eight Gell-Mann flavor matrices and $G_{S}$, $G_{V}$ and $K$ are the coupling constants. In the NJL model, there is no mediator, and the quark-quark interaction is a direct four-fermi contact interaction. The EoS for SQM derived from the Lagrangian (Eq. (4)) can be found in Lopes and Menezes (2021); Masuda et al. (2013). We consider the Hatsuda-Kunihiro (HK) parameter set from Hatsuda and Kunihiro (1994), and two values of the vector to scalar coupling ratio ($G_{V}/G_{S}$) to describe the SQS configurations.
@@ II.2.3 Color-flavor-locked (CFL) model
In the framework of MIT bag model Chodos et al. (1974), it is already depicted in Lugones and Horvath (2002) that the ground state of hadronic matter can be strange CFL matter spanning a broad parameter space of the gap parameter $\Delta$ (which is the gap of the QCD Cooper pairs), the bag constant $B$ and the mass of the strange quark $m_{s}$. To begin with, for the unpaired quark matter (UQM) system, the thermodynamic potential is $\Omega_{free}$ to which the pairing is introduced in terms of $\Delta$ to invoke the CFL phase i.e,
$\Omega_{CFL}=\Omega_{free}-\frac{3}{\pi^{2}}\Delta^{2}\mu^{2}+B$ (5)
where, $\mu$ is the chemical potential. The choice of $\Delta=0$ reduces to the UQM case Alford and Reddy (2003). As equal number of quark flavor contribution is considered, the absence of electrons makes the net matter charge neutral Rajagopal and Wilczek (2001). In practice the choice of $B$, $\Delta$ and $m_{s}$ are not totally independent of each other and is decided by the stability condition of the star. This also suggests that the three flavors of the quarks forming Cooper pairs have same Fermi momentum (and number density) at $T=0$ Lugones and Horvath (2002); Vásquez Flores and Lugones (2017). The EoS of the CFL QM derived from the thermodynamic potential (Eq. (5)) can be found in Lugones and Horvath (2002). The parameter set, used in the present work for obtaining the SQS configurations with the CFL quark model, is tabulated in Tab. 4.
@@ III Hadron-quark phase transition and hybrid star
We achieve hadron-quark phase transition and consequently obtain HS configurations with the KIDS-A (softest) and KIDS-D (stiffest) to describe the hadronic phase while the v-Bag model is considered to account for the quark phase. We also considered the other quark models like the NJL and CFL models described in Sec. II.2.2 and II.2.3, respectively to obtain HS configurations. However, we found that both NJL and CFL quark models along with the KIDS hadronic model yield unstable HSs. This result is also consistent with Lopes et al. (2022); Contrera et al. (2017); Alaverdyan (2022), where the NJL quark model along with the other hadronic models also give rise to unstable HS configurations. Similarly, in Kumar et al. (2024b) it can be seen that the CFL quark model combined with the other hadronic models, via Maxwell construction, forms unstable HS configurations. The unstable region of the HSs corresponds to the region where $dM/d\varepsilon_{c}<0$, where $\varepsilon_{c}$ is the central energy density of the star. Therefore, in the present work we consider the configurations of HSs with the KIDS-A, KIDS-D hadronic models and the v-Bag quark model. We present the corresponding results in the context of phase transition and HSs. In case of HSs, it is not mandatory to satisfy the Bodmer-Witten conjecture and the stability conditions in terms of the binding energy per baryon. Also we showed in Sen et al. (2024) that the values of $G_{V}$ and $B$ may be constrained in the light of the various astrophysical constraints on the structural properties of the compact stars. Therefore, for obtaining the HS configurations, we fix $X_{V}$ = 0.4, $B^{1/4}$ = 155 MeV and vary the values of $G_{V}$.
@@ IV Structural properties and oscillation modes
Using the EoS for various composition like HNS matter, SQM, and HS matter, we estimate the global properties like the gravitational mass ($M$) and the radius ($R$) of the haronic star, SQSs and HSs. The metric for a static spherically symmetric star is given as
@@
By solving the TOV equations for all possible values of $\varepsilon_{c}$, we obtain the mass $M$ and radius $R$ of the star. The mass function $M(r)=r(1-e^{-2\lambda(r)})/2$ is obtained by satisfying Eq. (10).
We calculate the non-radial oscillations of the hadronic star, SQSs and HSs using the Cowling approximations, which is well formulated in Sotani et al. (2011). Once we solve the TOV Eqs. (8 - 10), the different oscillation mode frequencies can be obtained by solving two more following coupled differential equations along with the corresponding coefficients obtained in terms of the solutions of the TOV equations:
$\frac{dW(r)}{dr}$ $=$ $\frac{d\varepsilon(r)}{dP(r)}\Bigg{[}\omega^{2}r^{2}e^{\lambda(r)-2\Phi(r)}V(r)+\frac{d\Phi(r)}{dr}W(r)\Bigg{]}-l(l+1)e^{\lambda(r)}V(r),$ (11)
@@
where, $A$ is an arbitrary constant. At the surface ($r=R$) of the star another boundary condition for $W(r)$ and $V(r)$ is given as
@@
$V_{+}=\frac{e^{2\Phi}}{\omega^{2}R_{g}^{2}}\left\{\frac{\varepsilon_{-}+P}{\varepsilon_{+}+P}\bigg{[}\omega^{2}R_{g}^{2}e^{-2\Phi}V_{-}+e^{-\Lambda}\frac{d\Phi}{dr}W_{-}\bigg{]}-e^{-\Lambda}\frac{d\Phi}{dr}W_{+}\right\}.$ (16)
@@ V.1 Hadronic Neutron Stars
In this section we present the structural and consequently the oscillation properties of the HNSs in Fig. 1. The non-relativistic KIDS model yields softer EoS compared to the RMF models. This is reflected in Fig. 1a, where we find that the maximum mass of the KIDS-A,B,C and D models is less than that of the RMF DD2 and GM1 models. However, the KIDS-A,B,C and D models satisfy the mass and corresponding radius of the most massive pulsar PSR J0740+6620 Fonseca et al. (2021); Miller et al. (2021); Riley et al. (2021). On the other hand, the radius of the neutron star having a given mass, predicted by the KIDS model is quite less compared to that by both DD2 and GM1 models, making the neutron stars more compact in the framework of the KIDS model. This helps all the four KIDS models satisfy not only the constraints from GW170817 Abbott et al. (2018), NICER experiment for PSR J0030+0451 Riley et al. (2019); Miller et al. (2019) but also HESS J1731-347 V. Doroshenko, V. Suleimanov, G. Pühlhofer, and A. Santangelo (2022); whereas, the last constraint is not satisfied by the DD2 and GM1 models. In Fig. 1b we study the variation of dimensionless tidal deformability with mass of the HNSs. It can be seen that all the KIDS models A, B, C, and D satisfy the constraint on $\Lambda_{1.4}$ from GW170817 Abbott et al. (2018) while on the contrary this constraint is not satisfied by the two RMF models considered in this work. Although the constraint on $\Lambda_{1.4}$ from GW190814 Abbott et al. (2020) is still speculative because the exact nature of the secondary object related to GW190814 emission is still unknown, this constraint is satisfied by the KIDS-C, and the DD2 and GM1 models.
@@
In Figs. 1c and 1d we portray the variation of $f$ and $p$ mode oscillation frequencies, respectively, of the HNSs with respect to the mass of the stars. The two RMF models, being stiffer than the KIDS model, have lower $f$ mode oscillation frequency compared to all the KIDS models. It is also clear from these two figures that the $R_{1.4}$ shows excellent correlation with $f_{f_{1.4}}$ but not with $f_{{p_{1}}_{1.4}}$. A careful look at Fig. 1a shows that $R_{1.4}$ for the different hadronic models increases in the trend KIDS-B $<$ KIDS-D $<$ KIDS-A $<$ KIDS-C $<$ DD2 $<$ GM1 while Fig. 1c shows that $f_{f_{1.4}}$ for the different hadronic models decreases by following the exact trend. Thus $f_{f_{1.4}}$ follows an inverse correlation with $R_{1.4}$ as seen in Kunjipurayil et al. (2022). In the present work the range of $R_{1.4}$ for HNSs is (12.09 - 13.33) km with corresponding range of $f_{f_{1.4}}$ as (2.02 - 2.3). On the other hand, we do not find any well correlation between $R_{1.4}$ and $f_{{p_{1}}_{1.4}}$ as we see from Fig. 1d that $f_{{p_{1}}_{1.4}}$ for the different hadronic models decreases by following the trend KIDS-D $>$ KIDS-C $>$ KIDS-B $>$ DD2 $>$ KIDS-A $>$ GM1. Comparing Figs. 1c and 1d, it can be said that the $p$ mode frequency is more affected by the model uncertainties than the $f$ mode. In other words, for any fixed mass of HNS, the difference in $f_{f}$ is quite less compared to that for $f_{p_{1}}$ for the different hadronic models. In case of $f$ mode, the difference between $f_{f_{1.4}}^{max}$ (KIDS-B) and $f_{f_{1.4}}^{min}$ (GM1) is 0.28 kHz while for $p_{1}$ mode the difference is 1.1 kHz. However, it is interesting to note from Fig. 1d that $f_{{p_{1}}_{1.4}}$ is well correlated with $L$. $f_{{p_{1}}_{1.4}}$ follows the exact reverse trend compared to that of $L$ : GM1 $>$ KIDS-A $>$ KIDS-B,C $>$ DD2 $>$ KIDS-D, following Tab. LABEL:Tab:KIDS. Since KIDS-B and C have same value of $L$, the $f_{{p_{1}}_{1.4}}$ for these two models are very close as seen from Fig. 1d.
@@
We now seek the percentage of linear correlation between the nuclear matter parameters, structural, and oscillation properties of HNSs by calculating the Pearson correlation coefficients. We display the corresponding results in Fig. 2. From the results of HNSs of masses 1.4 $M_{\odot}$ and 2.01 $M_{\odot}$ shown in Figs. 2a and 2b, respectively, we find that irrespective of the mass of HNSs, both the $f$ and $p$ mode frequencies are negatively correlated to all the nuclear matter properties like $K_{0}$, $J$, $L$ and $K_{\rm sym}$. Comparing the results for mass 1.4 $M_{\odot}$ and 2.01 $M_{\odot}$ we notice that among the various nuclear matter properties, $f$ mode frequency is most correlated to $K_{\rm{sym}}$ while $p$ mode frequency shows maximum correlation with $L$, irrespective of mass of HNSs. In both cases, $f$ mode frequency is also moderately correlated to $K_{0}$ and $L$. $f_{p}$ and $J$ are moderately correlated, whereas, the correlation between $f_{f}$ and $J$ is poor. The correlations between the $f$ and $p$ mode frequencies with the other nuclear matter parameters remain inconclusive as their magnitudes show huge fluctuations with mass of the HNSs. Coming to the structural properties like $\Lambda$ and $R$, we find that $f$ and $p$ mode frequencies show negative correlation with them. $f_{f}$ shows strongest correlation with both $\Lambda$ and $R$, irrespective of mass of HNSs while for $f_{p}$ the correlation with $\Lambda$ and $R$ is less compared to that of $f_{f}$ and also diminishes with decreasing mass of HNSs.

### v2

@@ II.1.1 KIDS functional
In addition to the rules for the expansion of the energy, KIDS formalism also assumes rules for fitting the model parameters. Equation (2) represents the energy in homogeneous infinite nuclear matter. Three $\alpha_{i}$ and four $\beta_{i}$ are adjusted to the input nuclear matter data. To apply the theory to finite nuclei, we transform the energy to Skyrme-type contact interactions so that it is easily implemented in the Hartree-Fock codes. Details for the steps toward finite nuclei are illustrated in Refs. Gil et al. (2017a, b); Papakonstantinou and Hyun (2023). In the Skyrme-type force, two terms describing the density gradient and the spin-orbit coupling are added to the nuclear matter functional. Two new parameters in the Skyrme-type force are fitted to reproduce selected nuclear data. Since the fitting to nuclear data is done on top of the parameters determined from the nuclear matter data, nuclear matter EoS can be treated independently of the nuclear properties.
@@
Following the rules described above, KIDS-A, B, C and D models are determined to satisfy the neutron star properties constrained by modern astronomical observations. The density dependence of the symmetry energy plays a crucial role in determining the nuclear matter EoS at densities below and above the saturation density. With the purpose to reduce the uncertainties in the density dependence of the symmetry energy, incompressibility of the symmetric matter $K_{0}$, and the symmetry energy parameters $J$, $L$ and $K_{\rm sym}$ are constrained to reproduce the nuclear data and neutron star observation simultaneously. The detailed process is explained in Gil and Hyun (2021). Table LABEL:Tab:KIDS tabulates the values of $K_{0}$, $J$, $L$ and $K_{\rm sym}$ for the KIDS-A, B, C and D models.
@@ II.1.2 RMF models
The nucleons interact via the scalar $\sigma$, vector $\omega$ and iso-vector $\rho$ mesons. The vacuum expectation values of the meson fields ($\sigma_{0}$, $\omega_{0}$ and $\rho_{03}$) in RMF approximation and the EoS obtained from the Lagrangian (Eq. (3)) can be found in Xia et al. (2022). For the GM1 model the mesons have density-independent couplings $g_{\sigma}$, $g_{\omega}$ and $g_{\rho}$ with the nucleons and $g_{2}$ and $g_{3}$ are the higher order scalar field coefficients while $c_{3}$ is the higher order vector field coefficient. These non-linear self couplings are effectively considered in order to account for the in-medium effects. On the other hand, in the DD2 model Typel et al. (2010) $g_{2}$=$g_{3}$=$c_{3}$=0 and the in-medium effects are treated with the density-dependent couplings $g_{i}(\rho)$ (where, $i=\sigma,\omega,\rho$) following the Typel-Wolter ansatz Lu et al. (2011). The values of the density-independent couplings for GM1 model and density-dependent couplings for DD2 model along with the other parameters of the two models can be found in the respective references and also in Guha and Sen (2024). The saturation properties like the saturation density $\rho_{0}$, symmetry energy $J$, slope $L$, nuclear incompressibility $K_{0}$, skewness coefficient $S_{0}$, and the curvature parameter $K_{\rm{sym}}$ of the nuclear symmetry energy, as obtained for the two RMF models considered in this present work, can be found in the respective references and also in Xia et al. (2022). For comparison with the KIDS model, we also display the values of $K_{0}$, $J$, $L$ and $K_{\rm sym}$ for the DD2 and GM1 models in Table LABEL:Tab:KIDS.
@@ II.2.1 Vector MIT bag (v-Bag) model
We consider that the quark matter is composed of the $u$, $d$ and $s$ quarks and the electrons. The masses of the quarks are $m_{u}$ = 5 MeV, $m_{d}$ = 7 MeV and $m_{s}$ = 95 MeV. Based on the MIT bag model framework Chodos et al. (1974) , in the v-Bag model the repulsion between the quarks is mediated by a vector meson of mass $m_{V}$ = 783 MeV Lopes et al. (2022); Kumar et al. (2022); Laskos-Patkos et al. (2024); Kumar et al. (2023c). The Lagrangian for the v-Bag model is given by Eq. 4.
@@
$V_{\mu}$ denotes the vector field, $g_{qqV}$ is the quark-vector meson coupling constant, $B$ is the bag constant, $V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$ and the Heaviside function $\Theta$=1 inside the bag. For the vector coupling constant, we assume $g_{uuV}=g_{ddV}$ and $X_{V}=g_{ssV}/g_{uuV}$ = 0.4. In case of QSs, the parameter $G_{V}=(g_{uuV}/m_{V})^{2}$ and its corresponding value of $B$ are constrained by the Bodmer-Witten conjecture which is related to the stability of the star. For a particular value of $G_{V}$, $B_{max}$ are determined by the binding energy of the star at the surface $\varepsilon/\rho_{B}\leq$ 930 MeV while $B_{min}$ is determined with the 2-flavor QM Farhi and Jaffe (1984); Torres and Menezes (2013); Lopes et al. (2022). The EoS for SQM derived from the Lagrangian (Eq. (4)) can be found in Lopes et al. (2022). The parameter set, used in the present work for obtaining the SQS configurations with the v-Bag model, is tabulated in Tab. 2.
@@ II.2.2 Nambu-Jona-Lasino (NJL) model
${\cal L}_{vNJL}$ $=$ $\bar{\psi}_{f}[\gamma^{\mu}(i\partial_{\mu}-m_{f})]{\psi}_{f}$ (5)
$+$ $G_{S}\sum_{a=0}^{8}\left[(\bar{\psi}_{f}\lambda_{a}\psi)^{2}+(\bar{\psi}_{f}\gamma_{5}\lambda_{a}\psi)^{2}\right]$
-- $G_{V}(\bar{\psi}_{f}\gamma^{\mu}\psi)^{2}$
-- $K\left\{det[\bar{\psi}(1+\gamma_{5})\psi]+det[\bar{\psi}(1-\gamma_{5})\psi]\right\},$
where $m_{f}$ are the current quark masses, $\lambda_{a}$ are the eight Gell-Mann flavor matrices and $G_{S}$, $G_{V}$ and $K$ are the coupling constants. In the NJL model, there is no mediator and the quark-quark interaction is a direct four-Fermi contact interaction. The EoS for SQM derived from the Lagrangian (Eq. (5)) can be found in Lopes and Menezes (2021); Masuda et al. (2013). We consider the Hatsuda-Kunihiro (HK) parameter set from Hatsuda and Kunihiro (1994), and two values of the vector to scalar coupling ratio ($G_{V}/G_{S}$) to describe the SQS configurations.
@@ II.2.3 Color-flavor-locked (CFL) model
In the framework of the MIT bag model Chodos et al. (1974), it is already depicted in Lugones and Horvath (2002) that the ground state of hadronic matter can be strange CFL matter spanning a broad parameter space of the gap parameter $\Delta$ (which is the gap of the QCD Cooper pairs), the bag constant $B$ and the mass of the strange quark $m_{s}$. To begin with, for the unpaired quark matter (UQM) system, the thermodynamic potential is $\Omega_{free}$ to which the pairing is introduced in terms of $\Delta$ to invoke the CFL phase i.e,
$\Omega_{CFL}=\Omega_{free}-\frac{3}{\pi^{2}}\Delta^{2}\mu^{2}+B,$ (6)
where $\mu$ is the chemical potential. The choice of $\Delta=0$ reduces to the UQM case Alford and Reddy (2003). As the equal number of quark flavor contribution is considered, the absence of electrons makes the net matter charge neutral Rajagopal and Wilczek (2001). In practice, the choice of $B$, $\Delta$ and $m_{s}$ are not totally independent of each other and is decided by the stability condition of the star. This also suggests that the three flavors of the quarks forming Cooper pairs have same Fermi momentum (and number density) at $T=0$ Lugones and Horvath (2002); Vásquez Flores and Lugones (2017). The EoS of the CFL QM derived from the thermodynamic potential (Eq. (6)) can be found in Lugones and Horvath (2002). The parameter set, used in the present work for obtaining the SQS configurations with the CFL quark model, is tabulated in Tab. 4.
@@ III Hadron-quark phase transition and hybrid star
We achieve hadron-quark phase transition and consequently obtain HS configurations with the KIDS-A (softest) and KIDS-D (stiffest) to describe the hadronic phase while the v-Bag model is considered to account for the quark phase. We also considered the other quark models like the NJL and CFL models described in Sec. II.2.2 and II.2.3, respectively, to obtain HS configurations. However, we found that both NJL and CFL quark models along with the KIDS hadronic model yield unstable HSs. This result is also consistent with Lopes et al. (2022); Contrera et al. (2017); Alaverdyan (2022), where the NJL quark model along with the other hadronic models also gives rise to unstable HS configurations. Similarly, in Kumar et al. (2024b) it can be seen that the CFL quark model combined with the other hadronic models, via Maxwell construction, forms unstable HS configurations. The unstable region of the HSs corresponds to the region where $dM/d\varepsilon_{c}<0$, where $\varepsilon_{c}$ is the central energy density of the star. Therefore, in the present work we consider the configurations of HSs with the KIDS-A, KIDS-D hadronic models and the v-Bag quark model. We present the corresponding results in the context of phase transition and HSs. In case of HSs, it is not mandatory to satisfy the Bodmer-Witten conjecture and the stability conditions in terms of the binding energy per baryon. Also, we showed in Sen et al. (2024) that the values of $G_{V}$ and $B$ may be constrained in the light of the various astrophysical constraints on the structural properties of the compact stars. Therefore, for obtaining the HS configurations, we fix $X_{V}$ = 0.4, $B^{1/4}$ = 155 MeV and vary the values of $G_{V}$.
@@ IV Structural properties and oscillation modes
Using the EoS for various composition like HNS matter, SQM, and HS matter, we estimate the global properties such as the gravitational mass ($M$) and the radius ($R$) of the haronic star, SQSs and HSs. The metric for a static spherically symmetric star is given as
@@
By solving the TOV equations for all possible values of $\varepsilon_{c}$, we obtain the mass $M$ and the radius $R$ of the star. The mass function $M(r)=r(1-e^{-2\lambda(r)})/2$ is obtained by satisfying Eq. (11).
We calculate the non-radial oscillations of the hadronic star, SQSs and HSs using the Cowling approximations, which is well formulated in Sotani et al. (2011). Once we solve the TOV Eqs. (9 - 11), the different oscillation mode frequencies can be obtained by solving two more coupled differential equations along with the corresponding coefficients obtained in terms of the solutions of the TOV equations:
$\frac{dW(r)}{dr}$ $=$ $\frac{d\varepsilon(r)}{dP(r)}\Bigg{[}\omega^{2}r^{2}e^{\lambda(r)-2\Phi(r)}V(r)+\frac{d\Phi(r)}{dr}W(r)\Bigg{]}$ (12)
-- $l(l+1)e^{\lambda(r)}V(r),$
@@
where $A$ is an arbitrary constant. At the surface ($r=R$) of the star another boundary condition for $W(r)$ and $V(r)$ is given as
@@
$V_{+}=\frac{e^{2\Phi}}{\omega^{2}R_{g}^{2}}\Bigg{\{}\frac{\varepsilon_{-}+P}{\varepsilon_{+}+P}\bigg{[}\omega^{2}R_{g}^{2}e^{-2\Phi}V_{-}+e^{-\Lambda}\frac{d\Phi}{dr}W_{-}\bigg{]}$
$-e^{-\Lambda}\frac{d\Phi}{dr}W_{+}\Bigg{\}}.$ (17)
@@ V.1 Hadronic Neutron Stars
In this section we present the structural and consequently the oscillation properties of the HNSs in Fig. 1. The non-relativistic KIDS model yields softer EoS compared to the RMF models. This is reflected in Fig. 1a, where we find that the maximum mass of the KIDS-A,B,C and D models is less than that of the RMF DD2 and GM1 models. However, the KIDS-A,B,C and D models satisfy the mass and corresponding radius of the most massive pulsar PSR J0740+6620 Fonseca et al. (2021); Miller et al. (2021); Riley et al. (2021). On the other hand, the radius of the neutron star for a given mass predicted by the KIDS model, is quite less compared to that by both DD2 and GM1 models, making the neutron stars more compact in the framework of the KIDS model. This helps all the four KIDS models satisfy not only the constraints from GW170817 Abbott et al. (2018), NICER experiment for PSR J0030+0451 Riley et al. (2019); Miller et al. (2019) but also HESS J1731-347 V. Doroshenko, V. Suleimanov, G. Pühlhofer, and A. Santangelo (2022). The last constraint is not satisfied by the DD2 and GM1 models. In Fig. 1b we study the variation of dimensionless tidal deformability with the mass of the HNSs. It can be seen that all the KIDS models A, B, C, and D satisfy the constraint on $\Lambda_{1.4}$ from GW170817 Abbott et al. (2018). This constraint is not satisfied by the two RMF models considered in this work. Although the constraint on $\Lambda_{1.4}$ from GW190814 Abbott et al. (2020) is still speculative because the exact nature of the secondary object related to GW190814 emission is still unknown, this constraint is satisfied by the KIDS-C, and the DD2 and GM1 models.
@@
In Figs. 1c and 1d we portray the variation of $f$ and $p$ mode oscillation frequencies, respectively, of the HNSs with respect to the mass of the stars. The two RMF models, being stiffer than the KIDS model, have a lower $f$ mode oscillation frequency compared to all the KIDS models. It is also clear from these two figures that the $R_{1.4}$ shows excellent correlation with $f_{f_{1.4}}$ but not with $f_{{p_{1}}_{1.4}}$. A careful look at Fig. 1a shows that $R_{1.4}$ for the different hadronic models increases in the trend KIDS-B $<$ KIDS-D $<$ KIDS-A $<$ KIDS-C $<$ DD2 $<$ GM1 while Fig. 1c shows that $f_{f_{1.4}}$ for the different hadronic models decreases by following the exact trend. Thus, $f_{f_{1.4}}$ follows an inverse correlation with $R_{1.4}$ as seen in Kunjipurayil et al. (2022). In the present work the range of $R_{1.4}$ for HNSs is (12.09 - 13.33) km with corresponding range of $f_{f_{1.4}}$ as (2.02 - 2.3). On the other hand, we do not find any well correlation between $R_{1.4}$ and $f_{{p_{1}}_{1.4}}$ as we see from Fig. 1d that $f_{{p_{1}}_{1.4}}$ for the different hadronic models decreases by following the trend KIDS-D $>$ KIDS-C $>$ KIDS-B $>$ DD2 $>$ KIDS-A $>$ GM1. Comparing Figs. 1c and 1d, it can be said that the $p$ mode frequency is more affected by the model uncertainties than the $f$ mode. In other words, for any fixed mass of HNS, the difference in $f_{f}$ is quite less compared to that for $f_{p_{1}}$ for the different hadronic models. In case of $f$ mode, the difference between $f_{f_{1.4}}^{max}$ (KIDS-B) and $f_{f_{1.4}}^{min}$ (GM1) is 0.28 kHz while for $p_{1}$ mode the difference is 1.1 kHz. However, it is interesting to note from Fig. 1d that $f_{{p_{1}}_{1.4}}$ is well correlated with $L$. $f_{{p_{1}}_{1.4}}$ follows the exact reverse trend compared to that of $L$ : GM1 $>$ KIDS-A $>$ KIDS-B,C $>$ DD2 $>$ KIDS-D, following Tab. LABEL:Tab:KIDS. Since KIDS-B and C have the same value of $L$, the $f_{{p_{1}}_{1.4}}$ for these two models are very close, as seen from Fig. 1d.
@@
We now seek the percentage of linear correlation between the nuclear matter parameters, structural, and oscillation properties of HNSs by calculating the Pearson correlation coefficients. We display the corresponding results in Fig. 2. From the results of HNSs of masses 1.4 $M_{\odot}$ and 2.01 $M_{\odot}$ shown in Figs. 2a and 2b, respectively, we find that irrespective of the mass of HNSs, both the $f$ and the $p$ mode frequencies are negatively correlated to all the nuclear matter properties like $K_{0}$, $J$, $L$ and $K_{\rm sym}$. Comparing the results for mass 1.4 $M_{\odot}$ and 2.01 $M_{\odot}$ we notice that among the various nuclear matter properties, $f$ mode frequency is most correlated to $K_{\rm{sym}}$ while $p$ mode frequency shows maximum correlation with $L$, irrespective of mass of HNSs. In both cases, $f$ mode frequency is also moderately correlated to $K_{0}$ and $L$. $f_{p}$ and $J$ are moderately correlated, whereas, the correlation between $f_{f}$ and $J$ is poor. The correlations between the $f$ and $p$ mode frequencies with the other nuclear matter parameters remain inconclusive, as their magnitudes show huge fluctuations with the mass of the HNSs. Considering the structural properties like $\Lambda$ and $R$, we find that $f$ and $p$ mode frequencies show negative correlation with them. $f_{f}$ shows strongest correlation with both $\Lambda$ and $R$, irrespective of mass of HNSs while for $f_{p}$ the correlation with $\Lambda$ and $R$ is less compared to that of $f_{f}$ and also diminishes with decreasing mass of HNSs.